OpenStudy (anonymous):

Could someone help me with this problem ???????????????? A rectangular box is x feet long and x feet wide. The volume of the box is (4x^8 + 3x^6 ) cubic feet. What polynomial represents the height of the box in feet?

OpenStudy (anonymous):

The volume is\[length \cdot width \cdot height\] which is \[x \cdot x \cdot height = x^2 \cdot height\] You also know that te volume of the box is \[4x^8+3x^6\]So: \[x^2 \cdot height = 4x^8 +3x^6\] \[x^2 \cdot height = x^2(4x^6 +3x^4)\] \[height = 4x^6 +3x^4\]